A simple model of ammonia consists of a pyramid of hydrogen atoms at the vertices of the equilateral base and the nitrogen atom at the apex. The N-H distance is 0•10 nm and the angle between two N--H bonds is 108
Find the center of gravity of the molecule. (Atomic weights: H N 14;
1
Expert's answer
2020-11-18T14:51:57-0500

Put this molecule in a 3d coordinate system so that the midpoint between two hydrogen atoms is at 0, and the line that crosses these two hydrogens is lying on a Y-axis so that the third hydrogen is on X-axis. Nitrogen atom is in positive Z-direction. We know that the centre of gravity of an equilateral triangle of side "a" is


"x_c=\frac{\sqrt{3}}{6}a,"

The Y-coordinate for our disposition is


"y_c=0."

What left is to find "a" and "z_c".

But on the edge of any triangle HNH we have a right triangle with base "a\/2", hypothenuse of 0.10 nm, and angle of "108\u00b0\/2=54\u00b0." So, we can find "a":


"a=2\cdot0.10\cdot\text{sin}54°=0.16\text{ nm}.\\\space\\
x_c=\frac{\sqrt{3}}{6}a=\frac{\sqrt{3}}{6}0.16=0.046\text{ nm}."

Regarding the z-coordinate: consider this molecule as a dumbbell with 3 hydrogens on one side and one nitrogen on the other, the distance between the 'weights' is the height of the pyramid:


"h=\sqrt{0.1^2-\bigg(\frac{\sqrt3}{3}a\bigg)^2}=0.038\text{ nm}."

Thus, the z-coordinate of the centre of gravity of the molecule is

"z_c=\frac{z_{\text{H}_3}\cdot m_{\text{H}_3}+z_{\text{N}}\cdot m_{\text{N}}}{m_{\text{H}_3}+m_{\text{N}}}=\frac{0+0.038\cdot14}{1\cdot3+14}=0.032\text{ nm}."

Therefore, the center of gravity has the following coordinates:

"x_c=0.046\text{ nm},\\
y_c=0,\\
z_c=0.032\text{ nm}."